is f(x)=4/3-1/4x a polynomial? if yes to what degree? if not why?
Answer by sachi(548)

f(x)=4/3-1/4x
if the question is f(x)=4/3-(1/4)*x then it is polynomial of degree 1
if the question is f(x)=4/3-1/(4*x) then it is not a polynomial as the exponent of x is -1 not a whole no